OpenStudy (summersnow8):

Revise the recipe to make 150 mL of the medium: Triptone: 17.0 g Distilled water: 1L equation: weight given/volume given = weight desired/volume desired I understand how to convert weight.... but I am no sure about the distilled water @photon336

OpenStudy (summersnow8):

For example: \[\frac{ 17.0 g }{ 1000 mL } = \frac{ x }{ 150 mL }\] \[(1000 mL) x = (17.0 g)(150 mL) \] \[x = (17.0 g \times 150 mL) / 1000 mL\] x = 2.55 g

OpenStudy (photon336):

The distilled water is the solvent, what you're putting the triptone into. you've got 1000 mL of water and 17.0 grams of triptone, and you want to figure out how much triptone you would need to put into 135 mL of water to have the same concentration.

OpenStudy (photon336):

So, Based on what you did, I think you're right.